# Point Could Mamba: Point Cloud Learning via State Space Model

### v1

@@ Abstract
footnotetext: This work was performed when Tao Zhang was an Intern at Skywork AI.
In this work, for the first time, we demonstrate that Mamba-based point cloud methods can outperform point-based methods. Mamba exhibits strong global modeling capabilities and linear computational complexity, making it highly attractive for point cloud analysis. To enable more effective processing of 3-D point cloud data by Mamba, we propose a novel Consistent Traverse Serialization to convert point clouds into 1-D point sequences while ensuring that neighbouring points in the sequence are also spatially adjacent. Consistent Traverse Serialization yields six variants by permuting the order of x, y, and z coordinates, and the synergistic use of these variants aids Mamba in comprehensively observing point cloud data. Furthermore, to assist Mamba in handling point sequences with different orders more effectively, we introduce point prompts to inform Mamba of the sequence’s arrangement rules. Finally, we propose positional encoding based on spatial coordinate mapping to inject positional information into point cloud sequences better. Based on these improvements, we construct a point cloud network named Point Cloud Mamba, which combines local and global modeling. Point Cloud Mamba surpasses the SOTA point-based method PointNeXt and achieves new SOTA performance on the ScanObjectNN, ModelNet40, and ShapeNetPart datasets.
Figure 1: Several pipelines of point cloud modeling. (a) denotes point-based methods, including point-based methods, such as PointNet [32], PointNet++ [33], PointMLP [28], and PointNeXt [35]. (b) is transformer-based methods, including Point Transformer [61] and Point-MAE [30]. (c) represents Mamba-based methods, which offer advantages of global modeling and linear computational complexity.
@@ 1 Introduction
Point cloud analysis has become a popular topic in understanding 3D and has drawn attention from the research community. Unlike 2D image processing, point clouds are composed of unordered and irregular point sets, which makes it hard to apply the 2D image processing methods directly. Thus, recent deep-learning-based approaches have various representations, such as voxel-based and point-based methods. The representative works, Point-Net and Point-Net++, adopt the MLP-based design with deep hierarchical local priors, shown in Fig. 1 (a). After that, many research works focus on advanced local geometric modeling via convolution, graph modeling, or attention.
@@
Meanwhile, with the rapid progress of vision transformers, several works enhance the global modeling with transformer structure in point could, as shown in Fig. 1 (b). In addition, transformer architectures also work effectively in mask point pre-training, 3D segmentation, and in-context learning. However, the computation and memory costs are still huge. Recently, state space models [8, 10] (SSMs) have been proven to model long-range dependency in sequential data. In particular, Mamba [8] is proven effective as Transformer [44] for several challenging NLP tasks. After that, recent works explore SSMs in various vision tasks, including image representation learning, medical segmentation, and low-level vision tasks. One concurrent work, PointMamba, directly uses the Mamab layer to model the global context. However, there are still significant performance gaps between PointMamba and previous point-based methods.
@@
In this work, we ask an essential question: Can we achieve an efficient point cloud analysis architecture using Mamba and surpass the performance of point-based methods? In particular, we introduce the Point Cloud Mamba (PCM), a combining local and global modelling framework that outperforms the SOTA point-based method PointNeXt [35].
@@
PCM utilizes Mamba to model the global features of point clouds while maintaining linear computational complexity. To process 3-D point cloud data effectively by Mamba layers, we propose Consistent Traverse Serialization (CTS) to serialize point clouds into a 1-D point sequence while ensuring that neighbouring points in the sequence are also adjacent in space. Then, CTS can easily derive six variants by simply permuting the order of x, y, and z coordinates. Surprisingly, when these six variants of CTS are combined, Mamba layers can more effectively model point cloud features because different variants provide different perspectives of the point cloud. To facilitate Mamba in handling specific point sequences better, we introduce order prompts to provide Mamba with the arrangement rules of the current point sequence. Finally, we propose simple spatial coordinate mapping as positional embedding for points, more suitable for irregular point cloud data than RoPE [40] and learnable embedding.
@@
Thanks to the improvements above, we successfully introduced Mamba into Point Cloud analysis and obtained Point Cloud Mamba (PCM). PCM outperforms the SOTA point-based method PointNeXt on three datasets: ScanObjectNet [43], ModelNet40 [50], and ShapeNetPart [58]. On the ShapeNetPart part segmentation dataset, PCM-Tiny surpasses PointNeXt with a 0.4 OA improvement, using only 39% of the parameter of PointNeXt.
@@
In summary, we have the following contributions in this work: 1) We introduce Mamba into point cloud analysis and construct a combined local and global modeling framework named Point Cloud Mamba. 2) We propose consistent traverse serialization, order prompts, and positional encoding based on spatial coordinate mapping to assist Mamba in better handling point cloud data. 3) Point Cloud Mamba is the first Mamba-based method that works well in point cloud analysis. It outperforms the SOTA point-based method PointNeXt on ScanObjectNet, ModelNet40, and ShapeNetPart datasets.
@@ 2 Related Work
3D Point Clouds Classification. Recent works have used deep neural networks to process 3D point clouds. In particular, representative works, PointNet [32] and PointNet++ [33], are the pioneering point-based approaches to directly handle the point clouds using MLPs. Meanwhile, several works explore graph-based modeling to utilize 3D geometric topology. Then, several works [47, 42, 39, 21, 55, 15] explore the local geometric features via different kernel modeling. Moreover, several works [24, 56, 51, 36, 4, 14, 23, 53, 60, 28] explore other point cloud architecture designs, including MLPs and transformers. Several works also explore different pre-training methods [31, 59] or in-context abilities [6, 45] inspired by the NLP field. Recently, state space models [10, 8] have achieved significant progress. Compared with transformers, they have advantages in efficient global modeling. A concurrent work [22] explores such architecture in point clouds. However, there are still significant performance gaps, compared with previous point cloud methods. For the first time, our work shows that Mamba architecture can achieve comparable or even better results than transformer-based models in 3D point clouds.
3D Visual Transformers. With the rise of the transformer in 2D version [5, 20, 2], several works [16, 11, 38, 41, 17] also explore transformer architectures in the point cloud. Earlier works [11, 61] have focused on the point cloud process. PCT [11] performs global attention directly to each point, following the ViT [5]. However, it has memory consumption and computational complexity issues. Point Transformer [61] solves this issue by introducing local attention. Then, the updated versions [49, 48] explore the different architectures to improve performance and efficiency. Inspired by these studies, our works combine local point processing and a new traverse serialization strategy, which leads to better results than direct SSM traverse.
@@ 3 Method
Figure 2: The architecture of our proposed Point Cloud Mamba. PCM consists of four stages, each comprising a geometric affine module and several mamba layers. Point downsampling is performed between stages.
The SSM-based architecture, Mamba [8], is attractive for point cloud representation learning due to its global modeling capability and linear computational complexity. However, Mamba is designed for the causal modeling of 1-D sequences, making it difficult to directly apply it to the modeling of non-causal 3-D point cloud data, posing many challenges to be addressed.
This section explores how to effectively integrate Mamba into architectures based on local modeling to capture global features. We first review PointMLP [28], a straightforward local modeling architecture, in Sec. 3.1, and then review Mamba, a global modeling architecture with linear complexity, in Sec. 3.2. Next, in Sec. 3.3, we introduce how to combine Mamba with PointMLP to obtain a point cloud network based on Mamba architecture. Finally, we introduce improving this naive Mamba-based network to Point Cloud Mamba in Sec. 3.4. We propose consistent traverse serialization, order prompt, new positional embedding, and more reasonable architectures to assist Mamba in modelling point cloud data better.
3.1 PointMLP Formulation
PointMLP [28] models the representation of point clouds using simple MLP and expands the receptive field of each point through point cloud downsampling and local feature aggregation. The process can be described by Equation 1 and 2:
$f_{i}^{l+1}=\Phi_{pos}(\mathcal{A}(\Phi_{pre}(GAM(\{f_{i,j}^{l}\})),|j=1,...,K)),$ (1)
@@ 3.1 PointMLP Formulation
PointMLP enhances the features of point clouds by iteratively performing the above process and downsampling the point cloud between different stages to expand the receptive field. However, even though the GAM combined with the point cloud downsampling strategy can expand the receptive field of the point cloud, PointMLP still struggles to extend the receptive field to the entire global point cloud, thus lacking in global modeling of the point cloud.
3.2 Mamba Formulation
The state-space equation can describe a multi-input, multi-output continuous system where the current inputs and states jointly determine the change in the state space of this system:
@@ 3.2 Mamba Formulation
Inspired by equations 4 and 5, some SSM-based models have been proposed and demonstrated superiority in long sequence modeling. However, early SSM-based methods like S4 [9] assume linear time invariance (LTI), which limits their modeling capabilities. Recently, Mamba [8], a new SSM-based model that introduces time-varying system parameters, has been introduced. Specifically, $\Delta$, $\overline{\mathbf{A}}$, and $\overline{\mathbf{B}}$ are all functions of $x_{t}$. Mamba demonstrates long sequence modeling capabilities comparable to Transformer and achieves linear computational complexity during inference following equation 5.
However, equation 5 is difficult to compute in parallel. It can be expanded and implemented using global convolution to enhance the efficiency of training Mamba on GPUs:
@@ 3.3 A Naive Mamba-based Point Cloud Network
Using Mamba [8] to replace the MLP operator in PointMLP [28] to achieve global modeling capabilities is a natural idea. However, this still requires addressing two challenges: 1) Point clouds are 3-D data, so how can we transform them into 1-D sequences? 2) Mamba is designed for causal modeling, so how can Mamba handle non-causal point cloud data? To address these challenges, we first adopt the z-order [29] serialization method to flatten 3-D point cloud data into 1-D sequences, allowing point cloud data to be processed by Mamba. Secondly, inspired by [62] and [25], we use bidirectional Mamba to allow each point to obtain features from any other point. At this point, we have implemented a naive Mamba-based network for the point cloud:
@@ 3.4 Point Cloud Mamba
Serialization strategy. How to convert 3-D point cloud data into a 1-D sequence that Mamba can handle is crucial. We find that point cloud sequences arranged according to specific rules can be better processed by Mamba compared to disordered point cloud sequences. For example, randomly flattening point cloud data into a 1-D point sequence and feeding it into Mamba for modelling will significantly underperform compared to using an ordered point sequence (84.1 vs 86.7). Based on this, we propose the Consistent Traverse Serialization (CTS) strategy, which ensures that adjacent points in the sequence are also adjacent in spatial position. Figure 3 illustrates how CTS converts 3-D point clouds into a 1-D sequence.
@@
where $\{c_{1}^{s},c_{2}^{s},c_{3}^{s}\}$ and $\{c_{1}^{g},c_{2}^{g},c_{3}^{g}\}$ are the spatial and grid coordinates of the points, and $N$ is the grid numbers. We design an encoding function that, given the coordinates of two dimensions, maps the coordinates to a code. Sorting the sequence according to the code ensures that adjacent points are contiguous in space.
$Code\_func(n_{1},n_{2})=\left\{\begin{aligned} n_{2}*N+n_{1},\quad\quad&n_{2}\%2=0\\ (n_{2}+1)*N-n_{1},\quad&n_{2}\%2\neq 0\\ \end{aligned}\right.$ (9)
@@
Additionally, combining multiple serialization strategies can effectively assist Mamba in better modeling point cloud features, as shown in Table 9. Specifically, we adopt different serialization strategies for the inputs of different Mamba layers, allowing Mamba to perceive the point cloud more comprehensively, thus significantly surpassing a single serialization strategy alone.
@@
As shown in Figure 4, we assigned $N_{p}$ learnable embeddings as order prompts for each serialization order. Before being processed by the Mamba layer, the point cloud sequence has the corresponding order prompts added to both the beginning and end of the sequence. Considering that the input feature dimensions of the Mamba layers at different stages may vary, we also allocate a shared Linear layer for all Mamba layers within each stage to map the order prompts to the required channel size.
Positional embedding. In sequence modeling, positional encoding is crucial and widely applied in text and image patch sequence modeling. For example, RoPE [40] and learnable positional embedding are widely used in text sequence modeling and image patch sequence modeling, respectively. However, we find that these positional encoding methods are unsuitable for point clouds due to the sparsity and irregular shapes. The spatial distances between any two adjacent points in a point cloud sequence vary significantly, making it difficult to use the sequence positional information to represent the spatial gaps between adjacent points accurately. We employ a shared positional mapping function to address this challenge to map the point spatial coordinates to positional embeddings with the same channel size as the features.
$Emd_{pos}=Linear(\{c_{1}^{s},c_{2}^{s},c_{3}^{s}\})$ (11)
@@
Architecture Mamba layers Serialization Channels Order Prompts
PCM-Tiny {1, 1, 2, 2} {[xyz]-[xzy]-[yxz, yzx]-[zxy-zyx]} {192, 192, 384, 384} 6
PCM {1, 2, 2, 4} {[xyz]-[xzy, yxz]-[yzx, zxy]-[zyx, hilbert, z, z-trans]} {384, 384, 768, 768} 6

### v2

# Project page: https://github.com/zhang-tao-whu/PCM

@@ Abstract
Recently, state space models have exhibited strong global modeling capabilities and linear computational complexity in contrast to transformers. This research focuses on applying such architecture in point cloud analysis. In particular, for the first time, we demonstrate that Mamba-based point cloud methods can outperform previous methods based on transformer or multi-layer perceptrons (MLPs). To enable Mamba to process 3-D point cloud data more effectively, we propose a novel Consistent Traverse Serialization method to convert point clouds into 1-D point sequences while ensuring that neighboring points in the sequence are also spatially adjacent. Consistent Traverse Serialization yields six variants by permuting the order of x, y, and z coordinates, and the synergistic use of these variants aids Mamba in comprehensively observing point cloud data. Furthermore, to assist Mamba in handling point sequences with different orders more effectively, we introduce point prompts to inform Mamba of the sequence’s arrangement rules. Finally, we propose positional encoding based on spatial coordinate mapping to inject positional information into point cloud sequences better. Point Cloud Mamba surpasses the state-of-the-art (SOTA) point-based method PointNeXt and achieves new SOTA performance on the ScanObjectNN, ModelNet40, and ShapeNetPart datasets. Code and model are available at https://github.com/SkyworkAI/PointCloudMamba.
Keywords:
3D Point Cloud State space models Mamba
Figure 1: Several pipelines of point cloud modeling. (a) denotes point-based methods with only local perception, including point-based methods, such as PointNet [37], PointNet++ [38], PointMLP [34], and PointNeXt [40]. (b) is the transformer-based method with global perception but quadratic computational cost, including Point Transformer [74] and Point-MAE [36]. (c) represents Mamba-based methods, which offer advantages of global modeling and linear computational complexity.
@@ 1 Introduction
Point cloud analysis [37, 38, 26, 74, 40, 57, 40] has become a popular topic in 3D understanding and has drawn attention from the research community. Unlike 2D image processing, point clouds are composed of unordered and irregular point sets, making it difficult to apply the 2D image processing methods directly. Thus, recent deep-learning-based approaches [37, 38, 26, 57, 34, 40, 74, 14] propose using various methods, such as voxel-based and point-based, for point cloud representation. The representative works, Point-Net [37] and Point-Net++ [38], adopt the MLP-based design with deep hierarchical local priors, as shown in Fig. 1 (a). After that, many research works [34, 3, 47, 53, 74, 14, 59, 58] focus on advanced local geometric modeling via convolution, graph modeling, or attention.
@@
Meanwhile, with the rapid progress of vision transformers, several works [74, 14, 59, 58] enhance the global modeling with transformer structure in point could, as shown in Fig. 1 (b). In addition, transformer architectures also work effectively in mask point pre-training, 3D segmentation, and in-context learning. However, the computation and memory costs are still huge. Recently, state space models [12, 13] (SSMs) have been proven to model long-range dependency in sequential data. In particular, Mamba [12] is proven effective as Transformer [52] for several challenging NLP tasks. After that, recent works [76, 30, 24, 2, 33, 65, 69, 42, 15] explore SSMs in various vision tasks, including image representation learning, medical segmentation, and low-level vision tasks. One concurrent work, PointMamba, uses the Mamba layer to model the global context. However, there are still significant performance gaps between PointMamba and previous point-based methods.
@@
In this work, we ask an essential question: Can we design an efficient point cloud analysis architecture using Mamba and surpass the performance of point-based methods? In particular, we introduce the Point Cloud Mamba (PCM), a combining local and global modeling framework that outperforms the SOTA point-based method PointNeXt [40].
@@
PCM utilizes Mamba architecture to model the global features of point clouds while maintaining linear computational complexity. To process 3-D point cloud data effectively by Mamba layers, we propose a novel Consistent Traverse Serialization (CTS) method to serialize point clouds into a 1-D point sequence while ensuring that neighboring points in the sequence are also adjacent in space. Then, CTS can easily derive six variants by simply permuting the order of x, y, and z coordinates. Additionally, when these six variants of CTS are combined, Mamba layers can more effectively model point cloud features. This is because the different variants provide various perspectives of the point cloud, ensuring that spatially adjacent points are also adjacent in a serialized point sequence. To help Mamba handle specific point sequences better, we introduce order prompts to provide Mamba with the arrangement rules of the current point sequence. Finally, we propose simple spatial coordinate mapping as positional embedding for points, more suitable for irregular point cloud data than RoPE [48] and learnable embedding.
@@
Thanks to the above improvements, we successfully introduced Mamba into Point Cloud analysis and obtained Point Cloud Mamba (PCM). PCM outperforms the SOTA point-based method PointNeXt on three datasets: ScanObjectNet [51], ModelNet40 [62], and ShapeNetPart [71]. On the ShapeNetPart part segmentation dataset, PCM-Tiny surpasses PointNeXt with a 0.4 OA improvement, using only 39% of the parameter of PointNeXt.
@@
In summary, we have the following contributions: 1) We introduce Mamba into point cloud analysis and construct a combined local and global modeling framework named Point Cloud Mamba. 2) We propose consistent traverse serialization, order prompts, and positional encoding based on spatial coordinate mapping to assist Mamba in better handling point cloud data. 3) Point Cloud Mamba is the first Mamba-based method that works well in point cloud analysis. It outperforms the SOTA point-based method PointNeXt on ScanObjectNet, ModelNet40, and ShapeNetPart datasets.
@@ 2 Related Work
3D Point Cloud Classification. Recent works have used deep neural networks to process 3D point clouds. In particular, representative works, PointNet [37] and PointNet++ [38], are the pioneering point-based approaches to handle the point clouds using MLPs directly. Meanwhile, several works explore graph-based modeling to utilize 3D geometric topology. Then, several works [57, 50, 46, 26, 66, 20, 11, 7, 77] explore the local geometric features via different kernel modeling. Moreover, several works [29, 67, 63, 41, 6, 18, 28, 64, 73, 34, 23, 8, 68] explore other point cloud architecture designs, including MLPs and transformers. Several works also explore different pre-training methods [36, 72, 43, 44, 56, 17, 19, 61, 75, 70] or in-context abilities [10, 55, 60] inspired by the NLP field. Recently, state space models [13, 12] have achieved significant progress. Compared with transformers, they have advantages in efficient global modeling. A concurrent work [27] explores such architecture in point clouds. However, there are still significant performance gaps compared with previous point cloud methods. Our PCM shows that Mamba architecture can achieve comparable or even better results than transformer-based models in 3D point clouds.
3D Visual Transformers. With the rise of the transformer in 2D version [9, 25, 4], several works [21, 14, 45, 49, 22, 31, 54] also explore transformer architectures in the point cloud. Earlier works [14, 74] have focused on the point cloud process. PCT [14] performs global attention directly to each point, following the ViT [9]. However, it has memory consumption and computational complexity issues. Point Transformer [74] solves this issue by introducing local attention. Then, the updated versions [59, 58] explore the different architectures to improve performance and efficiency. Inspired by these studies, our works combine local point processing and a new traverse serialization strategy, which leads to better results than direct SSM traverse.
@@ 3 Method
Figure 2: The architecture of our proposed Point Cloud Mamba. PCM encoder consists of four stages, each comprising a geometric affine module and several mamba layers. Point downsampling is performed between stages. The decoder only consists of point interpolation, feature concatenation, and MLP.
The SSM-based architecture, Mamba [12], is attractive for point cloud representation learning due to its global modeling capability and linear computational complexity. However, Mamba is designed for the causal modeling of 1-D sequences, making it difficult to be directly applied to the modeling of non-causal 3-D point cloud data, posing many challenges to be addressed.
This section explores how to effectively integrate Mamba into architectures based on local modeling to capture global features. We first review PointMLP [34], a local modeling architecture, and Mamba, a global modeling architecture with linear complexity, in Sec. 3.1. Then, in Sec. 3.2, we introduce how to combine Mamba with PointMLP to obtain a point cloud network based on Mamba architecture. Finally, we introduce improving this naive Mamba-based network to Point Cloud Mamba in Sec. 3.3. We propose consistent traverse serialization, order prompt, new positional embedding, and more reasonable architectures to assist Mamba in better modeling point cloud data.
3.1 Preliminaries
PointMLP formulation. PointMLP [34] models the representation of point clouds using simple MLP and expands the receptive field of each point through point cloud downsampling and local feature aggregation. The process can be described by Equ. 1 and 2:
$f_{i}^{l+1}=\Phi_{2}(\mathcal{A}(\Phi_{1}(GAM(\{f_{i,j}^{l}\})),|j=1,...,K)),$ (1)
@@ 3.1 Preliminaries
Mamba formulation. The state-space equation can describe a multi-input, multi-output continuous system where the current inputs and states jointly determine the change in the state space of this system:
@@
Inspired by Equ. 4 and 5, Mamba [12], a new SSM-based model that introduces time-varying system parameters, has been introduced. Specifically, $\Delta$, $\overline{\mathbf{A}}$, and $\overline{\mathbf{B}}$ are all functions of $x_{t}$. Mamba demonstrates long sequence modeling capabilities comparable to Transformer and achieves linear computational complexity during inference following Equ. 5.
However, Equ. 5 is difficult to compute in parallel. It can be expanded and implemented using global convolution to enhance the efficiency of training Mamba on GPUs:
@@ 3.2 A Naive Mamba-based Point Cloud Network
Architecture. Using Mamba [12] to replace the MLP operator in PointMLP [34] to achieve global modeling capabilities is a natural idea. However, this still requires addressing two challenges: 1) Point clouds are 3-D data, so how can we transform them into 1-D sequences? 2) Mamba is designed for causal modeling, so how can Mamba handle non-causal point cloud data? To address these challenges, we first adopt the z-order [35] serialization method to flatten 3-D point cloud data into 1-D sequences, allowing point cloud data to be processed by Mamba. Secondly, inspired by [76] and [30], we use bidirectional Mamba to allow each point to obtain features from any other point. At this point, we have implemented a naive Mamba-based network for the point cloud:
@@
Shortcomings. However, the naive Mamba-based network still has some shortcomings, only achieving 84.1 OA on ScanObjectNN, significantly underperforming modern point-based methods like PointNeXt. There are some reasons: 1) Using a single serialization method to convert 3-D point cloud data into 1-D point sequences significantly loses the spatial relational information between points. Specifically, even if a point is adjacent to many others in space, it can only be adjacent to adjacent points in the point sequence. Therefore, employing multiple serialization methods in combination to alleviate this loss of spatial relational information is necessary. 2) The architecture design of the naive Mamba-based network follows the point-based method PointMLP, which may not be optimal or even reasonable for Mamba-based architecture. Exploring what benefits Mamba in modeling point clouds is important and necessary. Please refer to the Sec. 4.3 for a detailed discussion.
@@ 3.3 Point Cloud Mamba
Firstly, PCM provides various point cloud sequences through the consistent traverse serialization strategy and its variants, thereby preserving the relational information between points to the fullest extent by traversing multiple point sequences. Secondly, PCM introduces order prompts to assist Mamba layers in better handling point sequences generated by different serialization methods and strengthens the spatial position information of points through positional encoding. Finally, we have designed a more reasonable overall architecture.
@@
Serialization strategy. How to convert 3-D point cloud data into a 1-D sequence that Mamba can handle is crucial. We find that Mamba can better process point cloud sequences arranged according to specific rules than disordered point cloud sequences. For example, randomly flattening point cloud data into a 1-D point sequence and feeding it into Mamba for modeling will significantly underperform compared to using an ordered point sequence (84.1 vs. 86.7). Based on this, we propose the Consistent Traverse Serialization (CTS) strategy, which ensures that adjacent points in the sequence are also adjacent in spatial position. Fig. 3 illustrates how CTS converts 3-D point clouds into a 1-D sequence.
@@
where $\{c_{1}^{s},c_{2}^{s},c_{3}^{s}\}$ and $\{c_{1}^{g},c_{2}^{g},c_{3}^{g}\}$ are the spatial and grid coordinates of the points, and $N$ is the grid number. We design an encoding function that, given the coordinates of two dimensions, maps the coordinates to a code. Sorting the sequence according to the code ensures that adjacent points are contiguous in space.
$Code\_func(n_{1},n_{2})=\left\{\begin{aligned} n_{2}\times N+n_{1},\quad\quad&n_{2}\%2=0\\ (n_{2}+1)\times N-n_{1},\quad&n_{2}\%2\neq 0\\ \end{aligned}\right.$ (9)
@@
Additionally, combining multiple serialization strategies can effectively assist Mamba in better modeling point cloud features, as shown in Tab. 10. Specifically, we adopt different serialization strategies for the inputs of different Mamba layers, allowing Mamba to perceive the point cloud more comprehensively, thus significantly surpassing a single serialization strategy alone.
@@
As shown in Fig. 4, we assigned $N_{p}$ learnable embeddings as order prompts for each serialization order. Before being processed by the Mamba layer, the point cloud sequence has the corresponding order prompts added to both the beginning and end of the sequence. Considering that the input feature dimensions of the Mamba layers at different stages may vary, we also allocate a shared Linear layer for all Mamba layers within each stage to map the order prompts to the required channel size.
Positional embedding. In sequence modeling, positional encoding is crucial and widely applied in text and image patch sequence modeling. For example, RoPE [48] and learnable positional embedding are widely used in text sequence modeling and image patch sequence modeling, respectively. However, we find that these positional encoding methods are unsuitable for point clouds due to their sparsity and irregular shapes. The spatial distances between any two adjacent points in a point cloud sequence vary significantly, making it difficult to use the sequence positional information to represent the spatial gaps between adjacent points accurately. We employ a shared positional mapping function to address this challenge. It maps the point spatial coordinates to positional embeddings with the same channel size as the features.
$Emd_{pos}=Linear(\{c_{1}^{s},c_{2}^{s},c_{3}^{s}\})\vspace{-2mm}$ (11)
@@
Architecture PCM-Tiny PCM
Mamba layers {1, 1, 2, 2} {1, 2, 2, 4}
Serialization {[xyz]-[xzy]-[yxz, yzx]-[zxy-zyx]} {[xyz]-[xzy, yxz]-[yzx, zxy]-[zyx, Hilbert, z, z-trans]}
Channels {192, 192, 384, 384} {384, 384, 768, 768}
Order Prompts 6 6
Architecture settings. As shown in Fig. 2, our network architecture consists of four stages, each incorporating a geometric affine module and several Mamba layers. We employ a simple decoder without Mamba layers for point cloud segmentation. The decoder solely conducts point cloud interpolation, concatenation with multi-stage encoder features, and channel transformation through MLP. The number of Mamba layers, serialization strategies, channel sizes, and order prompt counts in each stage of PCM-Tiny are illustrated in Tab. 1. Furthermore, we obtain PCM by increasing the number of Mamba layers and channel sizes and adopting new serialization methods accordingly.
4 Experiments
4.1 Experiment Setup